OpenStudy (anonymous):

Simplify the following expression.. (x/x-4 + 6/x+4) divided by 3/2x^2-32

OpenStudy (anonymous):

(x(x+4)+6(x-4))(2x^2-32)/3(x-4)(x+4)=(x(x+4)+6(x-4))2(x^2-16)/3(x^2-16)= =2x(x+4)+6(x-4))/3